If a line falls on the points (15, 15) and (6, 11), what is its slope? Enter your answer as a fraction in lowest terms. Use a slash mark ( / ) as the fraction bar.

Answer by nerdybill(7384)

Given two points:
(15, 15) and (6, 11)
.
slope = (y2 - y1)/(x2- x1)
slope = (11 - 15)/(6 - 15)
slope = (-4)/(-9)
slope = 4/9